A straight line passes through the points (2; -3) and (4 - 4). The coordinate of the x intercept of this line is ?


1
Expert's answer
2022-02-20T16:57:45-0500

Solution:

Given points (2; -3) and (4, - 4).

Equation of line: "y-y_1=(\dfrac{y_2-y_1}{x_2-x1})(x-x_1)"

"\Rightarrow y+3=(\dfrac{-4+3}{4-2})(x-2)
\\\Rightarrow y+3=(\dfrac{-1}{2})(x-2)
\\\Rightarrow 2y+6=2-x"

For x-intercept, put y=0

"\Rightarrow 0+6=2-x
\\\Rightarrow x=2-6
\\\Rightarrow x=-4"

x-intercept is (-4, 0).